OpenStudy (anonymous):

(MEDAL+FAN) x-6=8 a. x=-2 b. x=2 c. x=-14 d. x=14 please help me and show me how to do it.

OpenStudy (anonymous):

8+6=14 14-6=8, so the answer is 14

OpenStudy (anonymous):

oh. that's it? never really thought it through like that. May I ask some more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

1/4a=3 a. a= 3/4 b. a= 3 c. a= 12 d. a=16

OpenStudy (anonymous):

add 6 to both sides and whatever you get is the value of x

OpenStudy (anonymous):

3/.25=12 .25*12=3, so the answer is 12

OpenStudy (anonymous):

okay. can i ask some more?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

wrong thing lol

OpenStudy (anonymous):

2/3x=6 a. x=9 b. x=4 c. x=18 d. x=12

OpenStudy (anonymous):

9

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@OrangeMaster i cant see my smart score?

OpenStudy (anonymous):

lol?

OpenStudy (anonymous):

Many questions.

OpenStudy (anonymous):

now i can see my smart score......

OpenStudy (anonymous):

I can't...

OpenStudy (anonymous):

o.o